Consider the arithmetic sequence; 2,7,12,17,22,.... what is the ordinal of the term 102?
Answer by stanbon(57203)

Consider the arithmetic sequence; 2,7,12,17,22,.... what is the ordinal of the term 102?
---------
The first term is a=2
The difference is d=7-2=5
The nth term is a(n)=a+(n-1)d
EQUATION:
102=2+(n-1)5
100=(n-1)5
n-1=20
n=21
The ordinal is 21
Cheers,
Stan H.